OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3. PLEASE HELP!

OpenStudy (anonymous):

\[f(x)=2x+1\]\[g(x)=-5x+2\]\[f(g(x))=2(-5x+2)+1=-10x+4+1=-10x+5\]\[f(g(3))=-10(3)+5=-30+5=-25\]

OpenStudy (anonymous):

thank you